OpenStudy (konradzuse):

derivative of (e^x)/(1+x^2)

OpenStudy (konradzuse):

(v'u-u'v)/v^2

OpenStudy (konradzuse):

(2e^x) - (e^x(1+x^2))/ (1+x^2)

OpenStudy (konradzuse):

woops I messed up...

OpenStudy (konradzuse):

\[\frac{e^x(1+x^2) - e^x(2x)}{(1+x^2)^2}\]

OpenStudy (anonymous):

\[\frac{d}{dx}(\frac{e^x}{1+x^2})\] Let u=the numerator and v=denominator \[\frac{u'v-v'u}{v^2}\] I think you reversed the numerator and denominator

OpenStudy (anonymous):

yea that should be it

OpenStudy (konradzuse):

it says ti simplify my answer now on wolfman it does give a simplifiction, but I'm not sure how to get there. http://www.wolframalpha.com/input/?i=derivative+of+%28e%5Ex%29%2F%281%2Bx%5E2%29&lk=4

OpenStudy (anonymous):

hell no!

OpenStudy (konradzuse):

(1+x^2)(1+x^2) I could cancel out one of the (1+x^2) on top? :P

OpenStudy (konradzuse):

so it could end up e^x-(e^x(2x))/(1-x^2)

OpenStudy (konradzuse):

grr :(

OpenStudy (anonymous):

in your answer, factor out -e^x in the numerator....

OpenStudy (anonymous):

** e^x not -e^x

OpenStudy (konradzuse):

e^x((1+x^2) -(2x))?

OpenStudy (anonymous):

yep.... writing it in standard form: x^2 - 2x + 1 which can be factored as (x-1)^2

OpenStudy (konradzuse):

So they aren't bracketed, I was confused with the power rule from before too, but satellite was saying the 2 are seperated, I guess I couldn't do much since it was multiplied...